# AMBER - Advanced SegFormer for Multi-Band Image Segmentation: an application to Hyperspectral Imaging

### v1

@@ 5.1 Classification results
6 Conclusions
In this paper, we presented AMBER, an advanced SegFormer designed and fine tuned for hyperspectral image segmentation. By incorporating three-dimensional convolutions and preserving spatial dimensions, AMBER effectively captures both local and global features in multi-band images. Our extensive experiments on the Indian Pines, Pavia University, and PRISMA datasets show that AMBER in most cases outperforms traditional CNN-based methods and other state-of-the-art models in terms of Overall Accuracy, Kappa coefficient, and Average Accuracy.
The results indicate that AMBER not only improves classification accuracy but also exhibits robust generalization capabilities across different types of hyperspectral images. This demonstrates the potential of transformer-based architectures in advancing the field of hyperspectral image analysis. Future work could explore further optimizations and applications of AMBER in various remote sensing tasks, potentially leveraging additional data augmentation techniques and transfer learning strategies to enhance performance even further. Additionally, AMBER could be applied to astronomical data to extend its utility and effectiveness in analyzing complex and diverse datasets.
[…]
*Acknowledgements This work has been funded by project code PIR01_00011 “IBISCo”, PON 2014-2020, for all three entities (INFN, UNINA and CNR) and by the PNRR CN1 - ISP_S2_AGRINTESA for UNINA.

### v2

@@ 5.1 Classification Results
6 Conclusion
In this paper, we presented AMBER, an advanced SegFormer designed and fine-tuned for multi-band image segmentation. By incorporating three-dimensional convolutions and preserving spatial dimensions, AMBER effectively captures both local and global features in multi-band images. Extensive experiments on the Salinas, Indian Pines, Pavia University, and PRISMA datasets demonstrate that AMBER, in most cases, outperforms traditional CNN-based methods in terms of Overall Accuracy, Kappa coefficient, and Average Accuracy.
The results indicate that AMBER significantly improves classification accuracy while exhibiting robust generalization capabilities across diverse hyperspectral datasets. Compared to traditional CNN methods like 3D-CNN and contextual deep CNN, AMBER’s transformer-based architecture gives it a clear advantage by effectively integrating spectral and spatial features. This aligns with findings from prior studies that emphasize the importance of capturing global context in hyperspectral image analysis. Additionally, the segmentation performance on the PRISMA dataset highlights AMBER’s adaptability to satellite-derived hyperspectral data, which are often characterized by high spectral variability and noise.
Despite these advancements, some challenges remain. For example, while AMBER achieved superior accuracy on most datasets, the performance on certain complex spectral regions suggests potential limitations in handling extreme spectral variations. This opens avenues for further research, particularly in enhancing the model’s robustness through advanced regularization techniques or adaptive attention mechanisms.
In a broader context, the adoption of transformer-based architectures like AMBER signals a paradigm shift in hyperspectral image analysis. The ability to process spatial and spectral information simultaneously makes these models well-suited for a variety of remote sensing applications, such as land-cover classification, environmental monitoring, and disaster assessment. Future research could extend the capabilities of AMBER by integrating domain-specific knowledge, leveraging transfer learning, and incorporating more diverse datasets.
Moreover, AMBER addresses the challenge of segmenting astronomical objects, such as galaxies, by effectively processing multi-band images like radio interferometric data. Its ability to handle complex, high-dimensional datasets makes it a promising tool for advancing segmentation tasks in astronomy. Simultaneously, AMBER can be employed to segment three-dimensional medical images, such as Magnetic Resonance Imaging (MRI) or Computed Tomography (CT) scans, demonstrating its versatility in tackling diverse segmentation challenges across multiple domains.
In summary, AMBER offers a robust and scalable approach to hyperspectral image segmentation, setting the stage for future innovations in the field. By building upon this work, researchers can further refine transformer-based methods to tackle emerging challenges and expand their applicability across diverse scientific domains.
Acknowledgements
This work has been funded by project code PIR01_00011 “IBISCo”, PON 2014-2020, for all three entities (INFN, UNINA and CNR) and by the PNRR CN1 - ISP_S2_AGRINTESA for UNINA.